OpenStudy (anonymous):

what is special about the exponential function f(x) = e^x and its derivative?

OpenStudy (campbell_st):

well what's the derivative..?

geerky42 (geerky42):

Ya, just take derivative of e^x, and you will see what is special about it.

OpenStudy (anonymous):

so is -e^x the derivative of e^x?

geerky42 (geerky42):

where did you get -e^x from, because it's not right? Do you have derivative table you can use?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is the derivative e^x then?

geerky42 (geerky42):

Yesh, \(\dfrac{d}{dx}e^x = e^x\) This is what special about e^x. you derivative it and you still have e^x

OpenStudy (anonymous):

than you :)

geerky42 (geerky42):

welcome